OpenStudy (anonymous):

If cos(x+pi/3)+sin(x-pi/3)=A then A=?

OpenStudy (anonymous):

convert sin x into cos by sin x= cos (pi/2 -x) and then cos A + cos B

OpenStudy (anonymous):

thank you, so would sin(x-pi/3)= cos(pi/2-x)?

OpenStudy (anonymous):

\[\Large \frac{1}{2} cosx-\frac{\sqrt{3}}{2} \sin x+ \frac{1}{2}sinx +\frac{\sqrt{3}}{2}cosx\]

OpenStudy (anonymous):

\[\cos(A+B) = cosAcosB-sinAsinB\]\[\sin(A-B) =sinAcosB-cosAsinB\]

OpenStudy (anonymous):

Thank you , I really appreciate the help

OpenStudy (anonymous):

\[cosx(\frac{1}{2})- sinx(\frac{\sqrt{3}}{2})... \] do what I need help said

OpenStudy (anonymous):

could you explain how you came up with the (1/2) (sqrt3/2), etc. ?

OpenStudy (anonymous):